Which of the following will exert greater osmotic pressure -0.1M NaCl or 0.1 M glucose ? Give reason for your answer
1
Expert's answer
2019-10-28T07:49:21-0400

"NaCl \to Na^++Cl^-"

"i(" Vant Hoff factor ")" for "NaCl=2"

while Glucose being an organic covalent compound ,so it does not dissociate in water.

Vant Hoff factor for glucose"=1"

Osmotic pressure for glucose"=iCRT=1\times 0.1 \times RT=0.1RT"

Osmotic pressure for "NaCl=iCRT=2\times 0.1\times RT=0.2RT=" Twice of that of glucose.

Hence osmotic pressure for sodium chloride is double of that of glucose